DIODE-PUMPED TERAHERTZ PHOTOMIXING SPECTROMETER

J. M. VAN ROOIJ^{a}, \underline{A. S. PINE} and R. D. SUENRAM, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899.

We previously demonstrated a high-resolution terahertz spectrometer^{b} by mixing dye lasers in a fast low-temperature-grown (LTG)-GaAs photoconductor with submicron interdigital electrodes feeding a broadband complementary spiral antenna which radiates the coherent terahertz beat through a sample gas cell to a composite Si bolometer. We have upgraded the system efficiency, range and sensitivity by replacing the dye lasers with near-infrared diode lasers, using lower capacitance photomixers and detecting the far-ir radiation with a faster response hot-electron InSb bolometer. The fixed-tuned laser is a single-frequency (λ ≈ 852 nm) distributed-Bragg-reflector (DBR) GaAlAs diode, self-locked to a confocal interferometer; and the tunable diode laser is a broadband quantum-well (828 ≤ λ ≤ 858 nm) single-mode Littman-Metcalf external-cavity oscillator and a tapered planar amplifier, with precision linear piezoelectric tuning of up to 60 GHz and complete spectral coverage from 0.1 to 2 THz. Sample spectra will be given.

aAlso, Department of Physics, Catholic University of Nijmegen, the Netherlands.

bA.S. Pine, R.D. Suenram, E.R. Brown and K.A. McIntosh, J. Mol. Spectrosc. 175, 37 (1996)